# Cartesian subgroups in graph products of groups

### v1

@@ 5.4. Admissible subsets
Consider the following partial ordering on $\prod_{i=1}^{m}G_{i}:$ ${\underline{g}}\leqslant{\underline{h}}$ if and only if ${\underline{g}}={\underline{h}}(I)$ for some $I\subset[m]$. Clearly, an element ${\underline{g}}\in Q$ is maximal if and only if ${\underline{g}}$ is a maximal element with respect to $\leqslant$. Also, $Q$ is admissible if and only if $Q$ is finite and is a lower set (i.e. ${\underline{g}}\in Q$ whenever ${\underline{g}}\leqslant{\underline{h}}\in Q$). Hence the following properties are clear.
@@
Denote $J:=\operatorname{supp}{\underline{g}}$ and $X=B_{Q\setminus\{{\underline{g}}\}}.$ In the view of Lemma 5.5, it is sufficient to prove that $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}=\bigcup_{j\in J}\{x_{j}=0\}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}.$
@@ 5.5. Loops defined by words
For each $k_{i},g_{i}\in G_{i}$, consider a standard path in the topological space $\operatorname{cone}G_{i}$ from the point $k_{i}$ to the point $k_{i}g_{i}$ along two straight segments, $k_{i}\leadsto 0\leadsto k_{i}g_{i}$.
@@
Here at each step $t=1,\dots,N$ we change the $i_{t}$-th coordinate from $k_{i_{t}}$ to $k_{i_{t}}g_{i_{t}}$ by going along a standard path, while all the other coordinates are fixed. Hence the path ends at the point $(h_{1},\dots,h_{m})\in(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$, where $h_{i}=\prod_{t:i_{t}=i}g_{i_{t}}\in G_{i}$.
@@
The element $g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in G_{1}\ast\dots\ast G_{m}\cong\pi_{1}(BG_{1}\vee\dots\vee BG_{m})$ is represented by the concatenation of loops which correspond to elements $g_{i}\in G_{i}\cong\pi_{1}(BG_{i})$. Since the isomorphism $\pi_{1}((B\underline{G},\ast)^{\mathcal{K}})\cong\underline{G}^{\mathcal{K}}$ from Proposition 2.5 is natural with respect to maps of simplicial complexes, the word $g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in\underline{G}^{\mathcal{K}}\cong\pi_{1}((B\underline{G},\ast)^{\mathcal{K}})$ is represented by the same loop. On the other hand, this loop is the composition of $\lambda:[0,1]\to(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ with the homotopy equivalence $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}\simeq(E\underline{G},\underline{G})^{\mathcal{K}}$ and the natural projection $(E\underline{G},\underline{G})^{\mathcal{K}}\to(B\underline{G},\ast)^{\mathcal{K}}$. This projection is injective on fundamental groups, so the claim follows. ∎
@@
where $\pi:\underline{G}^{\mathcal{K}}\to\prod_{i=1}^{m}G_{i}$ is the natural projection and $\sim$ is the homotopy modulo endpoints. Different choices of decomposition $x=g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in\underline{G}^{\mathcal{K}}$ correspond to homotopic paths: for example, each edge $\{i,j\}\in\mathcal{K}$ gives rise to the relation $g_{i}g_{j}=g_{j}g_{i}$ in $\underline{G}^{\mathcal{K}}$ (hence to uncertainty in the decomposition) and to disjoint union of squares in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ (hence to homotopies between corresponding paths). We do not prove this fact since we do not use it in full generality.
@@
To each standard element $L_{\underline{g}}(i,J)\in D$ we assign the loop in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ given in Definition 5.10 using the sequence from Definition 4.1 (e.g. $L_{\underline{g}}(3,\{1,3,4\})$ defines the same path as the sequence $g_{1}g_{3}g_{4}g_{3}^{-1}g_{4}^{-1}g_{1}^{-1}$). Hence each word $w=L_{\underline{g}}(i_{1},J_{1})^{\pm 1}\cdot\dots\cdot L_{\underline{g}}(i_{N},J_{N})^{\pm 1}$ on such elements defines a loop in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$. By the proposition above, this loop corresponds to $w\in\operatorname{Cart}(\underline{G},\mathcal{K})$ under the isomorphism $\pi_{1}(B,\underline{1})\cong\operatorname{Cart}(\underline{G},\mathcal{K})$.
For a subspace $X\subset(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$, we say that the word $w$ defines the path in $X$ if the corresponding path in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ lies in $X$. Hence if $L_{\underline{g}}(i,J)$ defines a path in $X$, then it can be considered as an element in $\pi_{1}(X,\underline{1})$.
@@ 5.6. Fundamental groups of subspaces corresponding to admissible subsets
Then the group $\pi_{1}(B_{Q},\underline{1})$ is presented by $\sum_{{\underline{g}}\in Q}\widetilde{b}_{0}(\mathcal{K}_{\operatorname{supp}{\underline{g}}})$ generators, which correspond to words
@@
  • For each $I\subset[m]$ consider the point $v_{I}:={\underline{g}}([m]\setminus I)\in B_{Q}$ with the coordinates $(v_{I})_{j}=\begin{cases}1_{j},&j\in I\\ g_{j},&j\notin I\end{cases}$ and the path
@@
  • For each $i\in J$ consider the path $s_{i}:a_{\max(J)}\leadsto v_{\varnothing}\leadsto a_{i}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ from $a_{\max(J)}$ to $a_{i}$. (Here $s_{i}$ the concatenation of standard paths along the segments $[0_{\max(J)},g_{\max(J)}]$ and $[g_{i},0_{i}]$).
@@
  • For each edge $\{i,j\}\in\mathcal{K}_{J}$, consider the point $u_{ij}\in{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ with the coordinates $(u_{ij})_{k}:=\begin{cases}g_{k},&k\neq i,j\\ 0,&k\in\{i,j\}\end{cases}$ and the path $\ell(i,j):a_{i}\leadsto u_{ij}\leadsto a_{j}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$.
@@
This set of paths satisfies the conditions of Lemma 5.1 (here $PC=\Theta(J)\sqcup\{m^{\prime}\}$ and $x=\underline{1}$). Then Corollary 5.20 identify the resulting generators and relations in terms of words in $\operatorname{Cart}(\underline{G},\mathcal{K})$. This completes the proof of Theorem 5.14. ∎
@@ 5.7. Infinite case
Similarly, let $R\in\operatorname{Ker}(F(\Gamma)\to\pi_{1}(X))$ be a relation. It depends only on a finite set of generators $[\gamma_{1}],\dots,[\gamma_{N}]\in\Gamma.$ For every $j=1,\dots,N$ we have $[\gamma_{j}]\in\Gamma_{\alpha_{j}}\subset\pi_{1}(X_{\alpha_{j}})$ for some $\alpha_{j}.$ The relation $R$ corresponds to a homotopy between a loop in $X$ and the trivial map. The image of homotopy is an image of unit square, so it is compact and is contained in some $V_{\alpha_{0}}.$ For some $\alpha$, we have $V_{\alpha_{0}},V_{\alpha_{1}},\dots,V_{\alpha_{N}}\subset V_{\alpha}.$
It follows that the word $R$ represents is a relation between $[\gamma_{1}],\dots,[\gamma_{N}]$ in $\pi_{1}(X_{\alpha}),$ hence $R$ belongs to the normal closure of the subset $R_{\alpha}\subset F(\Gamma_{\alpha}).$ It follows that $R$ belongs to the normal closure of $\bigcup_{\alpha}R_{\alpha}\subset F(\bigcup_{\alpha}\Gamma_{\alpha}).$
@@ 6.2. Number of relations in Cartesian subgroups of graph products
they disagree if and only if $\operatorname{rank}\pi_{\operatorname{ab}}<\operatorname{rank}\pi.$ Often these ranks are equal. However, the ranks can differ even if all the groups $\Pi_{1}(\mathcal{K}_{J})$ are abelian: for example,
$21=\operatorname{rank}\mathbb{Z}_{6}=\operatorname{rank}(\mathbb{Z}_{2}\oplus\mathbb{Z}_{3})<\operatorname{rank}(\mathbb{Z}_{2}\ast\mathbb{Z}_{3})=$ 2 .
@@
Although $\operatorname{rank}\pi$ and $\operatorname{rank}\pi_{\operatorname{ab}}$ can differ significantly, they agree when $m$ is small enough: flag simplicial complexes with small $m$ have simpler homotopical properties compared to arbitrary simplicial complexes. For example, one needs $n+1$ vertices to triangulate an $n$-sphere, but $2n$ vertices for a flag triangulation. Also, these ranks agree for surface triangulations:
@@
Indeed: for $|\mathcal{K}_{J}|$ a closed oriented surface, this is well known. Otherwise $|\mathcal{K}_{J}|$ is a proper compact subset of a closed surface with finitely generated fundamental group, hence $\pi_{1}(\mathcal{K}_{J})$ is free by [FZ05, Corollary 7]. (More elementary, one can prove that $\pi_{1}(\mathcal{K}_{J})$ is free by collapsing, in the sense of simply homotopy theory, $\mathcal{K}_{J}$ onto its 1-skeleton.) ∎
Remark 6.4.
Due to Proposition 2.5 and [Dav08, Theorem 10.6.1] (see also [Cai17, Theorem 2.3]), $\mathcal{R}_{\mathcal{K}}$ is an aspherical $3$-manifold if and only if $\mathcal{K}$ is a flag triangulation of $S^{2}.$ In this case we know exact number of relations in a “minimal presentation” of the 3-manifold group $\mathrm{RC}^{\prime}_{\mathcal{K}}=\pi_{1}(\mathcal{R}_{\mathcal{K}})$. Note that $H_{1}(\mathcal{R}_{\mathcal{K}};\mathbb{Z})\simeq\bigoplus_{J\subset[m]}{\widetilde{H}}_{0}(\mathcal{K}_{J};\mathbb{Z})$ is free abelian, so the Poincaré duality gives $H_{1}(\mathcal{R}_{\mathcal{K}};\mathbb{Z})^{*}\simeq H_{2}(\mathcal{R}_{\mathcal{K}};\mathbb{Z}).$ Hence our presentation is balanced: the number of generators equals the number of relations. Existence of a balanced presentation is predicted by [Eps61, Theorem 2.5]. Theorem 4.11 provides a minimal balanced presentation of $\mathrm{RC}^{\prime}_{\mathcal{K}}$ in this case. It would be interesting to check if geometric methods of Grujić [Gru22] give a similar presentation.
@@ 6.3. Finite coverings and quotients of real moment-angle complexes
The group {±$\{\pm 1\}^{m}$ acts naturally on the real moment-angle complex $\mathcal{R}_{\mathcal{K}}=([-1,1],\{-1,1\})^{\mathcal{K}}$. Quotients $\mathcal{R}_{\mathcal{K}}/H$ by freely acting subgroups $H\subset\{\pm 1\}^{m}$ are known as real toric spaces. In this way one can construct manifolds which admit interesting geometric structures [WY21, Ero22]; the most important examples are the small covers, the real analogues of quasitoric manifolds [DJ91]. Fundamental groups of real toric spaces $\pi_{1}(\mathcal{R}_{\mathcal{K}}/H)$ are thus intermediate subgroups between right-angled Coxeter groups and their commutator subgroups, $\pi_{1}(\mathcal{R}_{\mathcal{K}}/H)\cong\operatorname{Ker}(\mathrm{RC}_{\mathcal{K}}\to\{\pm 1\}^{m}/H)\supset\mathrm{RC}_{\mathcal{K}}^{\prime}$. Classifying spaces of similar subgroups in the right-angled Artin groups were considered in [GILV18].
@@
Study the fundamental groups of real toric spaces using the methods of this paper.

### v2

@@ 5.4. Admissible subsets
Consider the following partial ordering on $\prod_{i=1}^{m}G_{i}:$ ${\underline{g}}\leqslant{\underline{h}}$ if and only if ${\underline{g}}={\underline{h}}(I)$ for some $I\subset[m]$. Clearly, $Q$ is admissible if and only if $Q$ is finite and is a lower set (i.e. ${\underline{h}}\in Q$ whenever ${\underline{h}}\leqslant{\underline{g}}$ for some ${\underline{g}}\in Q$). Also, ${\underline{g}}\in Q$ is maximal if and only if ${\underline{g}}$ is a maximal element of $Q$ with respect to $\leqslant$. Hence the following properties are clear.
@@
Denote $J:=\operatorname{supp}{\underline{g}}$ and $X:=B_{Q\setminus\{{\underline{g}}\}}.$ In view of Lemma 5.5, it is sufficient to prove that $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}=\bigcup_{j\in J}\{x_{j}=0\}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}.$
@@ 5.5. Loops defined by words
For each $k_{i},g_{i}\in G_{i}$, consider the standard path in the topological space $\operatorname{cone}G_{i}$ from the point $k_{i}$ to the point $k_{i}g_{i}$ along two straight segments, $k_{i}\leadsto 0\leadsto k_{i}g_{i}$.
@@
Here at each step $t=1,\dots,N$ we change the $i_{t}$-th coordinate from $k_{i_{t}}$ to $k_{i_{t}}g_{i_{t}}$ by following the standard path, while all the other coordinates are fixed. Hence the path ends at the point $(h_{1},\dots,h_{m})\in(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$, where $h_{i}=\prod_{t:i_{t}=i}g_{i_{t}}\in G_{i}$.
@@
The element $g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in G_{1}\ast\dots\ast G_{m}\cong\pi_{1}(BG_{1}\vee\dots\vee BG_{m})$ is represented by the concatenation of loops which correspond to the elements $g_{i}\in G_{i}\cong\pi_{1}(BG_{i})$. Since the isomorphism $\pi_{1}((B\underline{G},\ast)^{\mathcal{K}})\cong\underline{G}^{\mathcal{K}}$ from Proposition 2.5 is natural with respect to maps of simplicial complexes, the word $g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in\underline{G}^{\mathcal{K}}\cong\pi_{1}((B\underline{G},\ast)^{\mathcal{K}})$ is represented by the same loop. On the other hand, this loop is the composition of $\lambda:[0,1]\to(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ with the homotopy equivalence $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}\simeq(E\underline{G},\underline{G})^{\mathcal{K}}$ and the natural map $(E\underline{G},\underline{G})^{\mathcal{K}}\to(B\underline{G},\ast)^{\mathcal{K}}$. This map is injective on the fundamental groups, so the claim follows. ∎
@@
where $\pi:\underline{G}^{\mathcal{K}}\to\prod_{i=1}^{m}G_{i}$ is the natural projection and $\sim$ is the homotopy modulo endpoints. Different factorizations $x=g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in\underline{G}^{\mathcal{K}}$ correspond to homotopic paths: for example, each edge $\{i,j\}\in\mathcal{K}$ gives rise to the relation $g_{i}g_{j}=g_{j}g_{i}$ in $\underline{G}^{\mathcal{K}}$ (hence to ambiguity in the factorization) and to disjoint union of squares in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ (hence to homotopies between the corresponding paths). We do not prove this fact since we do not use it in full generality.
@@
To each standard element $L_{\underline{g}}(i,J)\in D$ we assign the loop in the space $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ as in Definition 5.10 using the sequence from Definition 4.1 (e.g. $L_{\underline{g}}(3,\{1,3,4\})$ is considered as the sequence $g_{1}g_{3}g_{4}g_{3}^{-1}g_{4}^{-1}g_{1}^{-1}$). Hence each word $w=L_{\underline{g}}(i_{1},J_{1})^{\pm 1}\cdot\dots\cdot L_{\underline{g}}(i_{N},J_{N})^{\pm 1}$ on the standard elements defines a loop in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$. By Proposition 5.11, this loop represents the element $w\in\operatorname{Cart}(\underline{G},\mathcal{K})$ under the isomorphism $\pi_{1}((\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}},\underline{1})\cong\operatorname{Cart}(\underline{G},\mathcal{K})$.
We say that such a word $w$ defines a loop in a subspace $X\subset(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ if the corresponding loop in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ lies in $X$. Hence if $L_{\underline{g}}(i,J)$ defines a loop in $X$, then it represents an element in $\pi_{1}(X,\underline{1})$, which is mapped into $L_{\underline{g}}(i,J)$ under the homomorphism $\pi_{1}(X,\underline{1})\to\pi_{1}((\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}},\underline{1})\cong\operatorname{Cart}(\underline{G},\mathcal{K})$.
@@ 5.6. Fundamental groups of subspaces corresponding to admissible subsets
Then the group $\pi_{1}(B_{Q},\underline{1})$ is presented by $\sum_{{\underline{g}}\in Q}\widetilde{b}_{0}(\mathcal{K}_{\operatorname{supp}{\underline{g}}})$ generators, which correspond to the words
@@
  • For each $I\subset[m]$ consider the point $v_{I}:={\underline{g}}([m]\setminus I)\in B_{Q}$ with coordinates $(v_{I})_{j}=\begin{cases}1_{j},&j\in I\\ g_{j},&j\notin I\end{cases}$ and the path
@@
  • For each $i\in J$ consider the path $s_{i}:a_{\max(J)}\leadsto v_{\varnothing}\leadsto a_{i}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ from $a_{\max(J)}$ to $a_{i}$. (Here $s_{i}$ the concatenation of the standard paths along the segments $[0_{\max(J)},g_{\max(J)}]$ and $[g_{i},0_{i}]$).
@@
  • For each edge $\{i,j\}\in\mathcal{K}_{J}$, consider the point $u_{ij}\in{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ with coordinates $(u_{ij})_{k}:=\begin{cases}g_{k},&k\neq i,j\\ 0,&k\in\{i,j\}\end{cases}$ and the path $\ell(i,j):a_{i}\leadsto u_{ij}\leadsto a_{j}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$.
@@
This set of paths satisfies the conditions of Lemma 5.1 (here $PC=\Theta(J)\sqcup\{m^{\prime}\}$ and $x=\underline{1}$). Then Corollary 5.20 expresses the resulting generators and relations in terms of words in $\operatorname{Cart}(\underline{G},\mathcal{K})$. This completes the proof of Theorem 5.14. ∎
@@ 5.7. Infinite case
Similarly, let $r\in\operatorname{Ker}(F(\Gamma)\to\pi_{1}(X))$ be a relation. It depends only on a finite set of generators $[\gamma_{1}],\dots,[\gamma_{N}]\in\Gamma.$ For every $j=1,\dots,N$ we have $[\gamma_{j}]\in\Gamma_{\alpha_{j}}\subset\pi_{1}(X_{\alpha_{j}})$ for some $\alpha_{j}.$ The relation $r$ corresponds to a homotopy between a loop in $X$ and the trivial map. The image of homotopy is an image of unit square, so it is compact and is contained in some $X_{\alpha_{0}}.$ For some $\alpha$, we have $X_{\alpha_{0}},X_{\alpha_{1}},\dots,X_{\alpha_{N}}\subset X_{\alpha}.$ Thus the word $r$ is a relation between the elements $[\gamma_{1}],\dots,[\gamma_{N}]$ of $\pi_{1}(X_{\alpha}),$ hence it belongs to the normal closure of the subset $R_{\alpha}\subset F(\Gamma_{\alpha}).$ It follows that $r$ belongs to the normal closure of $\bigcup_{\alpha}R_{\alpha}\subset F(\bigcup_{\alpha}\Gamma_{\alpha}).$
@@ 6.2. Number of relations in Cartesian subgroups of graph products
they disagree if and only if $\operatorname{rank}\pi_{\operatorname{ab}}<\operatorname{rank}\pi.$ This is possible even if all the groups $\Pi_{1}(\mathcal{K}_{J})$ are abelian: for example, 1=rank($21=\operatorname{rank}(\mathbb{Z}_{2}\oplus\mathbb{Z}_{3})<\operatorname{rank}(\mathbb{Z}_{2}\ast\mathbb{Z}_{3})=$ 2 .
@@
Although $\operatorname{rank}\pi$ and $\operatorname{rank}\pi_{\operatorname{ab}}$ can differ significantly, they agree when $m$ is small enough. Also, these ranks agree for surface triangulations:
@@
Indeed: for $|\mathcal{K}_{J}|$ a closed oriented surface, this is well known. Otherwise $|\mathcal{K}_{J}|$ is a proper compact subset of a closed surface with finitely generated fundamental group, hence $\pi_{1}(\mathcal{K}_{J})$ is free by [FZ05, Corollary 7]. (More elementary, one can prove that $\pi_{1}(\mathcal{K}_{J})$ is free by collapsing $\mathcal{K}_{J}$ onto its 1-skeleton.) ∎
Due to Proposition 2.5 and [Dav08, Theorem 10.6.1] (see also [Cai17, Theorem 2.3]), $\mathcal{R}_{\mathcal{K}}$ is an aspherical $3$-manifold if and only if $\mathcal{K}$ is a flag triangulation of $S^{2}.$ By [Eps61, Theorem 2.5], fundamental group of any aspherical $3$-manifold admits a balanced presentation, i.e. a presentation by $n$ generators modulo $n$ relations for some $n$. For $\mathcal{R}_{\mathcal{K}}$, our methods give a minimal balanced presentation.
Proposition 6.4.
Let $\mathcal{K}$ be a flag triangulation of $S^{2}$. Then Theorem 4.11 provides a presentation of the $3$-manifold group $\mathrm{RC}^{\prime}_{\mathcal{K}}$ by $N(\mathcal{K})$ generators modulo $N(\mathcal{K})$ relations, where $N(\mathcal{K})=\operatorname{rank}\mathrm{RC}^{\prime}_{\mathcal{K}}=\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})$.
Proof.
The group $H_{1}(\mathcal{R}_{\mathcal{K}};\mathbb{Z})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{0}(\mathcal{K}_{J};\mathbb{Z})\simeq\mathbb{Z}^{N(\mathcal{K})}$ is free abelian, so $\bigoplus_{J\subset[m]}{\widetilde{H}}_{1}(\mathcal{K}_{J};\mathbb{Z})\cong H_{2}(\mathcal{R}_{\mathcal{K}};\mathbb{Z})\simeq\mathbb{Z}^{N(\mathcal{K})}$ by the Poincaré duality. We have $\sum_{J\subset[m]}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})=N(\mathcal{K})$ by Proposition 6.3. Hence the presentation from Corollary 1.3 is minimal and balanced. ∎
It would be interesting to check if the geometric methods of Grujić [Gru22] give a similar presentation.
@@ 6.3. Finite coverings and quotients of real moment-angle complexes
The group {±$\{\pm 1\}^{m}$ acts naturally on the real moment-angle complex $\mathcal{R}_{\mathcal{K}}=([-1,1],\{-1,1\})^{\mathcal{K}}$. Quotients $\mathcal{R}_{\mathcal{K}}/H$ by freely acting subgroups $H\subset\{\pm 1\}^{m}$ are known as real partial quotients. In this way one can construct manifolds which admit interesting geometric structures [WY21, Ero22]; the most important examples are the small covers, the real analogues of quasitoric manifolds [DJ91]. The fundamental group of a real partial quotients $\pi_{1}(\mathcal{R}_{\mathcal{K}}/H)$ fits into the short exact sequence
$1\to\pi_{1}(\mathcal{R}_{\mathcal{K}}/H)\to\mathrm{RC}_{\mathcal{K}}\to\{\pm 1\}^{m}/H\to 1;$
in particular, $\mathrm{RC}^{\prime}_{\mathcal{K}}\subset\pi_{1}(\mathcal{R}_{\mathcal{K}}/H)\subset\mathrm{RC}_{\mathcal{K}}$. Classifying spaces of similar subgroups in the right-angled Artin groups were considered in [GILV18].
@@
Study the fundamental groups of real partial quotients using the methods of this paper.